Solve by substitution method
6x+5y=-2
x=12-7y
Can someone help me with this, thanks for your help

Answer by algebrahouse.com(1659)

6x + 5y = -2
x = 12 - 7y

6(12 - 7y) + 5y = -2 {substituted 12 - 7y, in for x, into top equation}
72 - 42y + 5y = -2 {used distributive property}
-37y + 72 = -2 {combined like terms}
-37y = -74 {subtracted 72 from both sides}
y = 2 {divided both sides by -37}

x = 12 - 7y {bottom original equation}
x = 12 - 7(2) {substituted 2, in for y, into 12 - 7y}
x = 12 - 14 {multiplied -7 by 2}
x = -2 {combined like terms}

x = -2 and y = 2
www.algebrahouse.com